OpenStudy (anonymous):

Suppose that a series sigma an has positive terms and its partial sums sn satisfy the inequality sn <= 1000 for all n. Explain why the sum of an must be convergent.

OpenStudy (rational):

Since all the terms are positive, the partial sums sequence is increasing. This is because adding a positive term always gives you a bigger number

OpenStudy (anonymous):

Ok.

OpenStudy (rational):

Also \(1000\) is an upper bound. So by monotone convergence theorem, the partial sums sequence converges. "The partial sums sequence converges" is same as saying "the series converges"

OpenStudy (anonymous):

Wait how is saying that when the sequence converges and so the series converges the same?

OpenStudy (anonymous):

How does that work in other words.

OpenStudy (rational):

sequence of "partial sums" converges is same as saying the series converges

OpenStudy (anonymous):

How? Could you give me an example please? :)

OpenStudy (rational):

lets look at a quick example maybe

OpenStudy (anonymous):

Hahaha, same idea!

OpenStudy (rational):

pick your favorite converging series

OpenStudy (anonymous):

Whoops. Oh yeah.

OpenStudy (anonymous):

Hahaha..ha.h.a.ha.

OpenStudy (rational):

pick some easy converging geometric series maybe

OpenStudy (anonymous):

1/n(n+1)?

OpenStudy (rational):

1/n(n+1) is NOT a series

OpenStudy (rational):

that looks like an algebraic expression to me

OpenStudy (rational):

a series must have "+" symbols or a \(\sum\) symbol

OpenStudy (anonymous):

5, -10/3, 20/9, -40/27 added together

OpenStudy (rational):

can you pick something simpler such that all terms are positive

OpenStudy (anonymous):

5, 10/3, 20/9, 40/27...? :)

OpenStudy (rational):

try picking a more simpler one

OpenStudy (rational):

how about 1/1, 1/2, 1/4, 1/8, 1/16, 1/32, ... ?

OpenStudy (anonymous):

Sure!

OpenStudy (rational):

thnks

OpenStudy (rational):

is that a sequence or series ?

OpenStudy (anonymous):

sequence.

OpenStudy (rational):

good, whats the first partial sum \(S(1)\) ?

OpenStudy (anonymous):

1

OpenStudy (rational):

whats the second partial sum \(S(2)\) ?

OpenStudy (anonymous):

S(2) = 3/2 S(3) = 7/4

OpenStudy (rational):

how about the first 6 parital sums ? \(S(1) = ?\) \(S(2) = ?\) \(S(3) = ?\) \(S(4) = ?\) \(S(5) = ?\) \(S(6) = ?\)

OpenStudy (perl):